OpenStudy (anonymous):

plz help:using the fact that 15=45-30, determine the exact value of sin 15 in simplest radical form

OpenStudy (anonymous):

@Hero plz help

hero (hero):

sin(a - b) = sin(a)cos(b) - cos(a)sin(b)

OpenStudy (anonymous):

is tht formula?

hero (hero):

sin(45 - 30) = sin(45)*cos(30) - cos(45)*sin(30)

hero (hero):

Yes that is the formula. All you have to do is just plug in the values and then simplify.

OpenStudy (anonymous):

what do u do with 15?

hero (hero):

Put it this way: sin (15) = sin(45 - 30) = sin(45)*cos(30) - cos(45)*sin(30)

OpenStudy (anonymous):

so this is wht I got:

OpenStudy (anonymous):

sin15=sin(15)=(raidcal2/2)(radical3/2)-(raidcal2/2)(radical1/2)

hero (hero):

Okay, so simplify it further

hero (hero):

sin(45 - 30) should stay as 45 - 30

OpenStudy (anonymous):

I shuld multiply radical2/2*radical3/2

hero (hero):

Yes, you should

OpenStudy (anonymous):

i got 1.224744871

hero (hero):

Not into decimals though

OpenStudy (anonymous):

raidcal6/2?

OpenStudy (anonymous):

it tht rit?

hero (hero):

Should be \(\dfrac{\sqrt{6}}{4}\) for that part

OpenStudy (anonymous):

ye thts what i meant

OpenStudy (anonymous):

so i got sin15=sin(15)=radical4/4?

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

@Hero

hero (hero):

Not exactly

hero (hero):

You should have ended up with \(\sin(15) = \dfrac{\sqrt{6} - \sqrt{2}}{4}\)

OpenStudy (anonymous):

oh thts my final answer? i cnt subtract radical6-radical2? bc the denominators are same

hero (hero):

How would you subtract \(\sqrt{6} - \sqrt{2}\)? Remember, you can't subtract if there is no property that allows you to.

hero (hero):

You can only do what the properties say that you can do.

hero (hero):

If you can find a radical or exponential property that allows you to subtract in that manner, then you can. Otherwise, you can't.

OpenStudy (anonymous):

oh got so thts my finasl answer? rit?

OpenStudy (anonymous):

@Hero

hero (hero):

Yes, that is the final answer

OpenStudy (anonymous):

thnks so much ur awesome! i appreciate ur time and help